Nick has two cans of white paint, one with a 2% yellow pigment and the other with a 5% yellow pigment. He wants to mix the two paints to get paint with a 4% yellow pigment. How much of the 5% yellow pigment paint should be mixed with 0.4 gallon of the 2% yellow pigment paint to get the desired paint?

Answer by stanbon(75887)

Nick has two cans of white paint, one with a 2% yellow pigment and the other with a 5% yellow pigment. He wants to mix the two paints to get paint with a 4% yellow pigment. How much of the 5% yellow pigment paint should be mixed with 0.4 gallon of the 2% yellow pigment paint to get the desired paint?
-----------------
Equation:
yellow + yellow = yellow
0.05x + 0.02*0.4 = 0.04(x+0.4)
----
Multiply thru by 100 to get:
5x + 2*0.4 = 4x+4*0.4
x = 2*0.4
x = 0.8 gallons (amt. of 5% paint needed in the mixture)
===================================
Cheers,
Stan H.